OpenStudy (caitlyn):

Evaluate 1/2a^-4b^2 for a=-2 and b=4 2 1/2 3 4 I got 128. I have a strange feeling that's wrong. So...yeah can someone please explain this?

hartnn (hartnn):

can you show how you got 128 ? then i'll tell you in which step, there was error.

OpenStudy (caitlyn):

I did (-2)^-4 and got 32 then 4^2 and got 16. Then I multiplied those two numbers then divided by 2.

hartnn (hartnn):

(-2)^-4 = 1/ (-2)^4 = 1/16 did you get how ?

hartnn (hartnn):

\[a^{-n} = 1/a^n\]

OpenStudy (caitlyn):

No. I don't understand how you got that. Whats the first step?

hartnn (hartnn):

\[(-2)^{-4} = \dfrac{1}{(-2)^4 }\\as, (a)^{-n} = \dfrac{1}{(a)^n }\] got this ?

OpenStudy (caitlyn):

Oh! yeah I forgot you had to do that fraction thing.

hartnn (hartnn):

\[(-2)^4=[(-2)^2]^2 = 4^2=16\] \[(-2)^2=(-2)(-2)=4\]

hartnn (hartnn):

so, now try again, you'll get it :)

OpenStudy (caitlyn):

Do you have to do the fraction thing for b?

hartnn (hartnn):

no, b^2 = 4^2 = 16

OpenStudy (caitlyn):

Ok when you do the fraction thing do you multiple the two numbers and then put 1 over fraction? 1/32 or 1/(-2)^-4

OpenStudy (caitlyn):

like 1/2x1/(-2)-4x16

OpenStudy (caitlyn):

i got 1/2

hartnn (hartnn):

1/2 is correct :)